A jet-liner and a propeller driven aircraft leave two cities 3000km apart at the same time and fly toward each other on a collision course. The flights are non-stop. If the propeller driven plane flies at 300km/h and the jet at 600km/h, how far will the jet fly before these planes crash?
Found 3 solutions by Alan3354, checkley77, nerdybill:
Answer by Alan3354(21603)

A jet-liner and a propeller driven aircraft leave two cities 3000km apart at the same time and fly toward each other on a collision course. The flights are non-stop. If the propeller driven plane flies at 300km/h and the jet at 600km/h, how far will the jet fly before these planes crash?
---------------------
Since they approach, their speeds are added, this is called the "rate of closure", the rate at which they converge.
300+600 = 900 km/hr closure speed
3000km/900 km/hr = 10/3 hours, or 3 hours 20 minutes until they collide.
The jet will fly 600*(10/3) km in that time, = 2000 km.
-----------------------------
BTW, airplanes, boats, ATC (Air Traffic Control) and the Coast Guard do NOT use kilometers and kph, they use knots and nautical miles. Schools and textbook writers are fond of metric (we all should be), but this is not used in aviation.
Also, it doesn't matter that one aircraft is a jet and one is a prop, or that the flights are non-stop.

Answer by checkley77(12569)

300x+600x=3000
900x=3000
x=3000/900
x=3.333 or 3 hours & 20 min. they will meet.
600*3.333=2,000 miles the jet will fly before crashing.

Answer by nerdybill(5405)

A jet-liner and a propeller driven aircraft leave two cities 3000km apart at the same time and fly toward each other on a collision course. The flights are non-stop. If the propeller driven plane flies at 300km/h and the jet at 600km/h, how far will the jet fly before these planes crash?
.
Applying the "distance formula".
.
Let t = time it takes to crash
.
300t + 600t = 3000
900t = 3000
t = 3000/900
t = 30/9
t = 3.33333 hours
or
3 hours 20 minutes